# How Memory Management Impacts LLM Agents: An Empirical Study of Experience-Following Behavior

### v1

@@ 4.3 Experience-Following Property
For each memory addition strategy, we measure both the input similarities and output similarities between each query and the memory records retrieved during its execution. The computation of both types of similarities for each query can be found in Appendix A.2. For each of the input similarities and the output similarities, we compute the cumulative average over the entire sequence of test query executions to demonstrate the long-term trend. As shown on the left of Figure 4 for AgentDriver, fixed-memory baseline produces both low input and output similarities, whereas addition-based methods exhibit higher output similarity as the input similarity grows. Similar patterns are observed for the other two agents (Figure 9 of Appendix B) and alternative LLM backbones such as state-of-the art GPT-4o and DeepSeek-V3 (Figure 15 of Appendix C.1). The exhibition of such a strong correlation between the input and output similarities is referred to as the experience-following property of memory-based LLM agents.
Figure 3: Left: Output similarity versus input similarity for AgentDriver over different addition strategies. Right: Average success rate versus input similarity of AgentDriver over different addition strategies.
Figure 4: Comparison of running performance between using the agent output as demonstrations and the error-free variant using ground-truth.
Intuitively, the experience-following property reveals the fact that LLM agents tend to follow the retrieved memory records more strictly when their queries are highly similar to the current task query. As the memory base expands by incorporating more experience from diverse use cases, the agent will retrieve records resembling highly similar past experiences more easily. Therefore, a careful selection strategy for memory addition can facilitate the self-improvement of the agent through the replication of correct executions of similar tasks. Conversely, indiscriminate memory addition will easily introduce incorrect executions for the agent to learn from, causing a self-degradation of its performance. To visualize this phenomenon, on the right of Figure 4, we show the relationship between the cumulative average input similarity and the success rate for different memory addition strategies for AgentDriver. We observe that strict selective addition fosters a high average performance with high input similarity, whereas methods that incorporate noisy entries into the memory yield significantly lower performance even with high input similarity. Similar trends can also be found for other agents as shown in Figure 10 in Appendix B.
@@ 4.4 Error Propagation in Agent Memory
While experience-following enables LLM agents to learn from high-quality examples when selective memory addition is adopted, the addition of erroneous memory records remains inevitable in many cases even with human inspection of the agent execution. For example, in AgentDriver, there is always a discrepancy between the optimal trajectory and the one predicted by the agent. When erroneous memory records are retrieved as demonstrations, these errors can influence the current task execution. If the current execution is then stored in the memory, the error may propagate to future tasks by affecting their execution. This error propagation poses significant challenges to memory management, causing a deviation between the agent’s performance and the optimal level achievable with perfect memory.
In Figure 4, we visualize error propagation on AgentDriver. For each addition strategy in Section 4.1, we compare it with a variant that uses the same retrieved examples for each task but replaces the LLM’s execution with the ground-truth trajectory, ensuring error-free [^2] While still suboptimal, the ground-truth trajectory resembles relatively “correct” task executions. retrieval. For each strategy, we observe an immediate gap in performance compared to its error-free variant. Moreover, as the execution continues, both add-all and coarse selective addition exacerbate such a performance gap, whereas strict selective addition, despite lagging initially, gradually approaches the ground-truth baseline and even surprisingly surpasses its performance after roughly 2000 executions. This behavior coincides with the high performance of AgentDriver using fixed, ground-truth memory without erroneous execution. However, when we add noise to the initial memory of AgentDriver (Appendix C.3), even with fixed memory, there is a rapid performance drop within 500 executions.
Interestingly, although with a much higher memory size, the performance of the error-free coarse addition strategy shows close performance to an error-free version of add-all. This observation naturally raises a question: Is retaining all memory records truly necessary to maintain a strong agent performance? This leads to the memory deletion strategies discussed in the sequel.
@@ 5 Deletion of Memory
Memory deletion is necessary for real systems, as the memory cannot grow indefinitely due to the practical constraint of storage. In this section, we evaluate three memory deletion approaches – periodical-based deletion, history-based deletion, and a deletion approach that comprehensively combines the two (Section 5.1). We present two key findings: (1) Strategic memory deletion can further enhance the agent’s performance while reducing the memory size: When used with selective addition, history-based memory deletion enhances the agent’s performance the most, while combined memory deletion reduces the memory size the most. (Section 5.2). (2) Misaligned experience replay: We identify misaligned experience replay, another challenge in memory management where some task-execution pairs are particularly difficult for the agent to benefit from experience-following. We show that the agent’s performance gain by history-based deletion can be attributed to the removal of the memory records associated with misaligned experience replay (Section 5.3).
@@ 5.1 Setup for Memory Deletion Experiments
Let $(q_{i},e_{i})$ be a query-execution pair where $1\leq i\leq N$ is the index. This pair will be removed from the memory bank if it satisfies a specific criterion $\phi$. We focus on three memory deletion strategies that can be deployed together with selective memory addition (either coarse (automatic) or strict (human) evaluation):
1) Periodical-based Deletion:
Prior research proposes to use a forgetting rate inspired by human cognition to assign deletion probabilities based on how frequently and how recently a memory record is retrieved zhong2024memorybank ; hou2024my . In this work, we adopt a simplified (thus more widely applicable) strategy using a fixed threshold to determine whether a memory record should be deleted based on its past retrieval frequency during a given period. Specifically, Let $\text{freq}_{t}(q_{i},e_{i})$ denote the total retrievals of $(q_{i},e_{i})$ at the current timestamp $t$, and $\text{freq}_{t^{\prime}}(q_{i},e_{i})$ be the retrievals at an earlier timestamp $t^{\prime}$. Define $\alpha$ as the target retrieval count within $[t^{\prime},t]$. A memory record will be deleted if $\phi_{period}(q_{i},e_{i},t,t^{\prime})=1$ where
$\phi_{period}(q_{i},e_{i},t,t^{\prime})\;=\;\mathds{1}\bigl{[}\text{freq}_{t}(q_{i},e_{i})\;-\;\text{freq}_{t^{\prime}}(q_{i},e_{i})\;\leq\;\alpha\bigr{]}.$
This approach keeps the memory size $M$ bounded by $M\leq\alpha(t^{\prime}-t)K$, where $K$ is the number of retrieved memory per execution.
2) History-Based Deletion:
We propose a history-based deletion strategy guided by the utility of stored memory records over time. Suppose $\Phi$ is a utility evaluator, which could be the same one used for selective addition. For any timestamp $t$, a record will be removed if a) it has been retrieved for at least $n$ times, and b) the average utility across all its past retrievals is below a prescribed threshold $\beta$:
$\phi_{history}(q_{i},e_{i},t)=\begin{cases}\mathds{1}\left[\frac{1}{\text{freq}_{t}(q_{i},e_{i})}\sum_{m=1}^{\text{freq}_{t}(q_{i},e_{i})}\Phi(q_{m},e_{m})\leq\beta\right],&\text{if }\text{freq}_{t}(q_{i},e_{i})>n,\\ 0,&\text{otherwise}.\end{cases}$
Note that we require a memory record to be retrieved at least $n$ times before being considered for removal to reduce the estimation bias for the average utility. Moreover, a memory record with high utility at timestamp $t$ may still be removed in the future as the task distribution evolves. We will discuss this scenario in Section 6.1.
3) Combined Deletion:
Periodical- and history-based methods can be applied together to jointly enhance the agent performance and reduce the memory size:
$\phi_{combine}(q_{i},e_{i},t,t^{\prime})\;=\;\phi_{period}(q_{i},e_{i},t,t^{\prime})\times\phi_{history}(q_{i},e_{i},t)$
@@ 5.2 Strategic Memory Deletion Improves the Agent Performance
Table 2: Performance of periodical-based deletion, history-based deletion, and the combined approach, when deployed with the two selective addition strategies (strict and coarse) and using the same evaluator for history-based and combined deletion. History-based deletion enhances the agent’s performance in most cases, while the combined method reduces the memory size the most.
@@
Strategy EHRAgents AgentDriver CIC-IoT Agent ACC. $\uparrow$ Mem Size $\downarrow$ SR. $\uparrow$ Mem Size $\downarrow$ ACC. $\uparrow$ Mem Size $\downarrow$ Coarse evaluator No del 31.35 1881 37.04 1161 60.67 1083 Period 30.30 379 36.39 426 59.58 163 History 32.04 1818 34.02 1019 53.08 1115 Combined 29.52 377 35.63 372 53.83 160 Strict evaluator No del 38.89 1012 50.94 1178 63.33 860 Period 39.04 302 50.94 467 61.83 148 History 42.65 784 51.81 846 67.67 755 Combined 42.88 248 49.81 323 63.50 82
@@
Table 2 summarizes the effects of three deletion strategies—periodical-based, history-based, and combined deletion—on both agent performance and memory size.
@@
We first observe that periodical-based deletion achieves substantial memory reduction with minimal performance degradation. Across most configurations, this method leads to absolute accuracy drops of less than 2%, indicating that addition-only memory designs often accumulate redundant entries.
History-based deletion and combined deletion show more variable results depending on the quality of the utility evaluator used during addition and deletion. Specifically, when a strict evaluator is employed, both strategies lead to notable performance improvements.
@@
To further understand these improvements, we extend the methodology used in Section 4.4 to construct an error-free memory baseline during agent execution. As presented in Appendix 16, history-based deletion with a strict evaluator can even outperform this error-free counterpart with a larger magnitude compared with no deletion. This result suggests that selectively retaining helpful experiences observed during execution can lead to better long-term performance by adapting more effectively to evolving task demands.
However, when the evaluator used for both addition and deletion is coarse or noisy, deletion methods can negatively affect agent performance. This highlights the challenge of using a coarse evaluator—its inherent errors can propagate, particularly when it drives the selection criteria for history-based and combined deletion. Indeed, because the coarse evaluator relies on an LLM with limited task-specific knowledge, it can misjudge execution quality and sometimes remove beneficial examples. In practice, this drawback can be mitigated by fine-tuning the evaluator on specific tasks to reduce errors.
@@
In summary, our findings show that with a reliable utility evaluator, history-based deletion can substantially improve agent performance. Additionally, the combined deletion approach offers a strong balance between maintaining performance and reducing memory size, making it a practical and efficient memory management strategy for long-term LLM agent deployment.
@@ 5.3 Misaligned Experience Replay
To explain the performance gains achieved by history-based deletion with a strict utility evaluator, we hypothesize that some memory records offer limited guidance during task execution. These records, which exhibit relative high input similarity to a task yet yield low output similarity when used as demonstrations, force the agent to rely more on its own reasoning—thus diminishing the benefits of experience following. We call this phenomenon misaligned experience replay to highlight the misalignment between the current task execution and the execution of its neighboring tasks retrieved from the memory bank as demonstrations. We attribute the performance gain of the history-based deletion to its successful removal of memory records associated with misaligned experience replay.
To visualize how history-based memory deletion handles misaligned experience replay, we divide a sequence of testing tasks into two groups: (1) deleted group: tasks for which the retrieved memory record exhibiting the highest output similarity has been deleted from the memory bank, ultimately and (2) remained group: the rest of the tasks with the top-output-similarity record retained in memory after all task executions. We then calculate the input and output similarity with the retrieved record that has the highest output similarity – we find clearly different patterns between the two groups.
Figure 5: Correlation between input similarity and output similarity over tasks leveraging deleted records and those leveraging retained records. History-based deletion consistently removes poor demonstrations with low output similarity, thereby improving long-term performance.
Considering the tasks on EHRAgent, for example, in Figure 5, tasks in the deleted group demonstrate that deleted records have much lower output similarity than retained records despite having comparable input similarity. The highlighted regions of the figure reveal that these deleted records lack effective experience-following – even when the input similarity is high, they do not lead to closely matched outputs, which degrades the long-term agent performance. Similar results for the other two agents can be found in Figures 11 and 12 in Appendix B.
@@ 6.1 Memory Management with Task distribution shift
Task distribution shift occurs when the predominant task type and the expected execution change over time. To simulate a task distribution shift, we construct a modified dataset from the original test sets of EHRAgent and AgentDriver. Specifically, we obtain the embedding vectors for all test queries using OpenAI text-embedding-3-large encoder for EHRAgent, and the original input vector computation method from AgentDriver. We then apply a Gaussian Mixture Model (GMM) to cluster the embeddings into three groups – each group will likely have a distinguished task distribution. Finally, we reorder the test queries based on their assigned cluster labels to process the three task groups sequentially. This procedure forces a pronounced shift in the task distribution.

### v2

@@ 3.3 Experience-Following Property
For each memory addition strategy, we measure both the input similarities and output similarities between each query and the memory records retrieved during its execution. The computation of both types of similarities for each query can be found in Appendix A. For each of the input similarities and the output similarities, we compute the cumulative average over the entire stream of test query executions to demonstrate the long-term trend. As shown in Figure 4 for RegAgent and AgentDriver, the fixed-memory baseline produces both low input and output similarities, whereas the addition-based methods exhibit higher output similarity as the input similarity grows. Similar patterns are observed for the other two agents (Figure 9 in Appendix B.1) and alternative LLM backbones such as state-of-the art GPT-4o and DeepSeek-V3 (Figure 12 in Appendix B.2). The exhibition of such a strong correlation between the input and output similarities is referred to as the experience-following property of memory-based LLM agents.
Intuitively, this property reflects that LLM agents tend to imitate retrieved experience more closely when their current queries are similar to past examples. As the memory expands to include a broader range of experiences, the agent is more likely to retrieve highly similar records for new queries. For instance, in RegAgent, the agent can make predictions either by directly mimicking the closest demonstration or by internal reasoning to guess the mapping rule based on demonstrations. When the memory size increases and demonstrations become more similar to the query, the agent shows a near-perfect correlation (Pearson $r\approx 1$) between input and output similarity, indicating strong demonstration following when input similarity is high. In contrast, under the fixed-memory setting, both input and output similarities remain consistently lower than those achieved with memory addition, suggesting that the agent still relies on reasoning rather than pure imitation, and thus achieves surprisingly high performance compared with those using a coarse evaluator with noisy memory. A similar pattern is also observed in AgentDriver. Therefore, a careful selection strategy with a stronger evaluator for memory addition can facilitate the self-improvement of the agent through the replication of correct executions of similar tasks. Conversely, indiscriminate memory addition will easily introduce incorrect executions for the agent to learn from, causing a self-degradation of its performance.
@@ 3.4 Error Propagation in Agent Memory
While experience-following enables LLM agents to learn from high-quality examples with a strong evaluator, the inclusion of erroneous memory records or seemingly well-performed trajectories with incorrect intermediate attempts remains inevitable in many cases, even under human inspection of the agent’s execution.
When erroneous and noisy memory records are retrieved as demonstrations, these errors can influence the current task execution. If the current execution is then stored in the memory, the error may propagate to future tasks by affecting their execution. This error propagation poses significant challenges to memory management, causing a deviation between the agent’s performance and the optimal level achievable with perfect memory.
In Figure 4, we visualize error propagation on RegAgent and AgentDriver. For each addition strategy in Section 3.1, we compare it with a variant that uses the same retrieved examples for each task but replaces the LLM’s execution with the ground-truth output, ensuring error-free [^2] For AgentDriver, while these trajectories can still be suboptimal, the ground-truth trajectory resembles relatively “correct” task executions. retrieval.
For both agents, we observe an immediate gap in performance compared to their error-free variant. Moreover, as the execution continues, both add-all and coarse selective addition exacerbate such a performance gap. However, for AgentDrievr with strict selective addition, despite lagging initially, it gradually approaches the ground-truth baseline and even surprisingly surpasses its performance after roughly 2000 executions. These results again highlight the importance of both the quantity and quality of memory, underscoring the necessity of carefully selecting and training trajectory evaluators for stable long-term agent execution.
@@ 4 Deletion of Memory
Memory deletion is often necessary for real systems, as the memory cannot grow indefinitely due to the practical constraint of storage. In this section, we evaluate three memory deletion approaches: the commonly adopted periodic deletion, the proposed history-based deletion guided by future execution utility, and a hybrid approach that combines both.
@@ 4.1 Setup for Memory Deletion Experiments
We follow Section 3.1 for the four agent setups we considered. For the deletion, let $(q_{i},e_{i})$ be a query-execution pair where $1\leq i\leq N$ is the index. This pair will be removed from the memory bank if it satisfies a specific criterion $\phi$. We focus on three memory deletion strategies that can be deployed together with selective memory addition (either coarse (automatic) or strict (human) evaluation):
1) Periodical-based Deletion: Prior research proposes to use a forgetting rate inspired by human cognition to assign deletion probabilities based on how frequently and how recently a memory record is retrieved Zhong et al. (2024); Hou et al. (2024); Wang et al. (2024b). In this work, we adopt a simplified (thus more widely applicable) strategy using a fixed threshold to determine whether a memory record should be deleted based on its past retrieval frequency during a given period. Specifically, Let $\text{fr}_{t}(q_{i},e_{i})$ denote the total retrievals of $(q_{i},e_{i})$ at the current timestamp $t$, and $\text{fr}_{t^{\prime}}(q_{i},e_{i})$ be the retrievals at an earlier timestamp $t^{\prime}$. Define $\alpha$ as the target retrieval count within $t^{\prime},t]$. A memory record will be deleted if $\phi_{\text{per}}(q_{i},e_{i},t,t^{\prime})=1$ where
$\phi_{\text{per}}(q_{i},e_{i},t,t^{\prime})=\mathbf{1}\!\left[\,\operatorname{fr}_{t}(q_{i},e_{i})-\operatorname{fr}_{t^{\prime}}(q_{i},e_{i})\leq\alpha\,\right].$
This approach keeps the memory size $M$ bounded by $M\leq\alpha(t-t^{\prime})K$, where $K$ is the number of retrieved memory per execution.
2) History-Based Deletion: Following the previous discussion on experience-following and error propagation, we hypothesize a correlation between the quality of experiences stored in memory and the downstream execution quality of future tasks when these experiences are retrieved. Thus, we propose a simple history-based deletion strategy guided by the utility of stored memory records over time. Specifically, suppose $\Phi$ is a utility evaluator, which could be the same one used for selective addition. For any timestamp $t$, a record will be removed if a) it has been retrieved for at least $n$ times, and b) the average utility across all its past retrievals is below a prescribed threshold $\beta$:
$\phi_{\text{hist}}(q_{i},e_{i},t)=\begin{cases}\delta(q_{i},e_{i},t),&\text{if }\text{fr}_{t}(q_{i},e_{i})>n,\\ 0,&\text{otherwise}.\end{cases}$
where
$\delta(q_{i},e_{i},t)=\mathbf{1}\!\left[\tfrac{1}{\operatorname{fr}_{t}(q_{i},e_{i})}\sum_{m=1}^{\operatorname{fr}_{t}(q_{i},e_{i})}\Phi(q_{m},e_{m})\;\leq\;\beta\right].$
Here, $\Phi(q_{m},e_{m})$ denotes the utility obtained from the $m$-th retrieval of the specific memory record $(q_{i},e_{i})$. Note that we require a memory record to be retrieved at least $n$ times before being considered for removal to reduce the estimation bias for the average utility.
3) Combined Deletion: Periodical- and history-based methods can be applied together to jointly enhance the agent’s performance and reduce the memory size:
$\phi_{\text{comb}}(q_{i},e_{i},t,t^{\prime})=\phi_{\text{per}}(q_{i},e_{i},t,t^{\prime})\,\lor\,\phi_{\text{hist}}(q_{i},e_{i},t).$
@@ 4.2 Strategic Memory Deletion Improves the Agent Performance
Table 2: Performance of periodical-based deletion, history-based deletion, and the combined approach, when deployed with the two selective addition strategies (strict and coarse with C1 evaluator) and using the same evaluator for history-based and combined deletion. The history-based approach in real agents often yields the best performance when paired with a strict evaluator among deletion approaches, while the combined method reduces the memory size the most.
@@
Strategy RegAgents EHRAgents AgentDriver CIC-IoT Agent SR. $\uparrow$ Mem Size $\downarrow$ ACC. $\uparrow$ Mem Size $\downarrow$ SR. $\uparrow$ Mem Size $\downarrow$ ACC. $\uparrow$ Mem Size $\downarrow$ Coarse evaluator (C1) No del 63.18 3511 25.91 1447 36.92 1161 74.00 1030 Period 60.88 1012 26.65 338 36.38 426 78.10 355 History 62.10 3205 33.55 1004 34.00 1019 73.70 952 Combined 59.32 951 31.47 279 35.62 372 68.80 352 Strict evaluator No del 70.95 2938 38.67 1012 51.00 1178 85.40 904 Period 67.65 949 38.59 302 50.94 467 80.80 310 History 69.80 2286 42.06 784 51.81 846 89.60 788 Combined 66.58 890 42.34 248 49.97 323 85.50 188
Figure 5: Performance comparison after applying history-based deletion with different evaluators.
@@
Table 2 summarizes the effects of three deletion strategies on both agent performance and memory size.
@@
We first observe that periodical-based deletion achieves substantial memory reduction with small performance degradation in general, indicating that addition-only memory designs often accumulate redundant entries.
History-based deletion and combined deletion show more variable results depending on the reliability of the utility evaluator used during addition and deletion. Specifically, when a strict evaluator is employed, history-based deletion leads to notable performance improvements with the non-synthetic agent we considered. In addition, for our synthetic RegAgent, which exhibits strong experience-following behavior and relies on a large memory capacity, we observe that history-based deletion causes only a small performance drop compared with other strategies, even after removing several hundred frequent executions.
@@
To further understand these improvements, we extend the methodology used in Section 3.4 to construct an error-free memory baseline during agent execution with AgentDriver and RegAgent. As presented in Appendix B.3, in AgentDriver, history-based deletion with a strict evaluator can even outperform this error-free counterpart with a larger magnitude compared with no deletion. This result suggests that selectively retaining experiences with high utility scores on downstream tasks can improve long-term performance.
However, when the evaluator used for both addition and deletion is coarse or noisy, history-based and combined deletion can influence the agent’s long-term behavior in different ways. As shown in Figure 5 and Figure 14 (Appendix B.4), the impact of history-based deletion varies notably across evaluators based on LLMs without fine-tuning. For instance, when using GPT-4o-mini as the history-based evaluator, the deletion strategy yields a clear performance gain on EhrAgent but leads to degraded performance on AgentDriver. In contrast, fine-tuned evaluators tend to produce more stable results, achieving performance comparable to their addition-only counterparts while maintaining smaller memory sizes.
@@
In summary, our findings show that with a reliable utility evaluator, history-based deletion can improve agent performance. Additionally, the combined deletion approach offers a strong balance between maintaining performance and reducing memory size, making it a practical and efficient memory management strategy for long-term LLM agent deployment.
@@ 4.3 Misaligned Experience Replay
Figure 6: Left: KDE curve over the absolute error of deleted and retained memory (with retrieved more than 5 times) with C1 evaluator using history-based deletion. Right: KDE curve over the absolute error with strict evaluator.
To explain the performance gains achieved by history-based deletion, we hypothesize that certain memory records provide limited or even harmful guidance during task execution. These records, although initially passing the evaluator’s quality filter, may still be misaligned with the objectives of the current task distribution. Such misalignment can arise from inconsistencies between their stored trajectories and the current execution context, or from errors introduced by the evaluator’s limitations. Conditioning on such demonstrations can therefore degrade the quality of task execution. We refer to this phenomenon as misaligned experience replay, emphasizing the misalignment between the current task and its retrieved demonstrations from the memory bank. We attribute the effectiveness of history-based deletion to its ability to remove these misaligned memory records.
Since inconsistencies between context and task distributions are often case-specific and difficult to capture directly, we instead use the error induced by the evaluator’s limitations for clearer illustration, as both factors jointly contribute to misaligned experience replay. Taking RegAgent as an example, the experience quality in terms of the error in the memory record can be directly observed through the discrepancy between its predicted and ground-truth value. As shown in Figure 6, we plot the kernel density estimation (KDE) curves for both deleted and retained records based on two different evaluators, respectively, revealing a clear quality gap: retained demonstrations exhibit lower error scores than deleted ones. Even under strict addition—where only data with error $\leq 1$ were initially stored—conditioning on noisier experiences (e.g., entries with error >0.5>0.5) can still propagate errors to subsequent tasks more easily, leading to degraded predictions and thus removed by history-based deletion. Consistent trends across other agents (Appendix B.5) further demonstrate that history-based deletion effectively identifies and removes noisy or misaligned memory records with various evaluators.
These findings suggest that experience quality, especially the intrinsic quality of experience itself, is tightly linked to downstream execution quality. By effectively removing those demonstrations with low future execution quality through an evaluator, we can mitigate misaligned experience replay.
@@ 5.1 Memory Management with Task distribution shift
Task distribution shift occurs when the predominant task type and the expected execution change over time. To simulate a task distribution shift, we construct a modified dataset from the original test sets of EHRAgent and AgentDriver by aggregating their input vectors and reordering them to alter the underlying task distribution. More details for setup can be found in Appendix A.5.